Ralph, treasurer for Petrova Imports, recently updated his firm's short-term cash forecast and discovered that the firm will suffer a cash shortage of $15 million over the next month (i.e., a 30 day period).  Ralph just learned from one of his investment bankers that commercial paper in the 30-day maturity range is in demand and that discount rates are roughly 3.7%.  The dealer's annual fee is 0.10% and the annual commitment fee on a backup line of credit is 0.25%. What is the effective borrowing cost on the commercial paper?
Answer #1

1. Face Value of commercial paper= $15 million

Issue price = Face value - (Discount rate * (Days/360) * Face Amount)

Issue price = $15 Million - 15 Million * 3.7% * 30 / 360

Issue price = 14,953,750

Interest paid = Face Value - Issue Price = 46250

Dealer fee= Annual Fee * Loan Amount * 30 / 360 = 0.1%*15000000*30/360 = $150000*30/360 = $1250

Commitment fee= Charge * Loan * 30 /360 0.25%*15,000,000*30/360= $ 3125

Total cost= 46250 + 1250 + 3125 = $50625

Effective rate= 50625 / 14,953,750 * 360/30

Effective rate= 4.06%